A battery of electric motive force of 24 volt internal resistance of 1 ohms connected to a resistor of 32 ohms.what is the current of the battery?


1
Expert's answer
2020-09-30T12:25:49-0400

solution

given data

electric motive force(E)=24V

resistance of battery(r)=1ohm

connected resistor(R)=32 ohm

total resistance (Req)=r+R

so

"R_{eq}=1+32\\=33\Omega"

then

by ohm law

"I=\frac{E}{R_{eq}}"


"I=\frac{24}{33}\\I=0.72A"

so current of battery is 0.72 A .